OpenStudy (anonymous):

Here's a trig question: cos 3x = 0.5 i think it starts at cos(2x + x) but i'm not sure

OpenStudy (tkhunny):

Is there a Domain over which we should be finding solutions? Maybe \(0 \le x \le 2\pi\)? Yu can certainly use that identity if you would like, but that may be the hardest way possible to solve the problem.

OpenStudy (anonymous):

the domain is 0 - 360 i'm pretty lost on this. another one i'm having trouble with is tan 2x = sq.root(3). if i now how to do these problems i can figure out the rest in this chapter, but i'm just having a really difficult time solving these kind of identities once it gets to the double/triple angles. keep in mind there may be multiple answers

OpenStudy (tkhunny):

There we go. That will help. If, in 0 <= x <= 360, we have cos(3x) = 0.5, we have immediately 3x = 60 and 3x = 300. You should know about 30-60-90 right triangles.

OpenStudy (anonymous):

okay i'll work from there. thank you